What is 3/20 (three twentienth) of 64?
Could someone please explain how to solve?

Answer by Alan3354(69443)

What is 3/20 (three twentienth) of 64?
--------
(3/20)*96 = 3*96/20
= 288/20
= 72/5
= 14.4